Find X if log 3 (x-4) +log 3 (x-2)=1 ............. ............

  • Find X if  log3(x-4) +log3(x-2)=1 .............    ............

Grade:11

2 Answers

Vikas TU
4 years ago
Dear student 
Log a + log b = log ab 
Log 3 ( x-4) (x-2) = 1 
3 = x^2 -6x + 8 
x^2 – 6x +5 = 0 
x = 2,3 
Hope this helps 
Arun
4 years ago
 
Log a + log b = log ab 
Log 3 ( x-4) (x-2) = 1 
3 = x^2 -6x + 8 
x^2 – 6x +5 = 0
hence x = 1 and 5
 
but 1 is not in its domain
hence rejected
 
hence only x = 5 is answer